OpenStudy (anonymous):

The lengths of the sides of a triangle are y,y+ 1, , and 7 centimeters. If the perimeter is 56 centimeters, what is the value of y?

OpenStudy (anonymous):

please show work...please HELP

OpenStudy (anonymous):

Recall, a triangle of sides A, B, C, has perimeter=A+B+C

OpenStudy (anonymous):

The sum of all sides is equal to 56. That's: \[y+y+1+7=56\] Can you solve this equation?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Fine. \[y+y+1+7=56 \implies 2y+8=56\] Subtract 8 from both side: \[2y=48 \implies y=24\]

OpenStudy (anonymous):

No need to be rude...not my fault i dont get it

OpenStudy (anonymous):

LOL.. I am not being rude :).. What didn't you get?

OpenStudy (anonymous):

nevermind...i get it thanks

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

What?